OpenStudy (anonymous):

3 sqrt symbol 49 over 3 sqrt symbol 2 how do you slimpfy the expression completely ?

OpenStudy (anonymous):

\[\frac{3\sqrt{49}}{3\sqrt{2}}\] First cancel a 3 from numerator and denominator:\[\frac{\sqrt{49}}{\sqrt{2}}\] Now simplify, given that the square root of 49 is 7: \[\frac{7}{\sqrt{2}}\] Now rationalize, multiplying numerator and denominator by the square root of 2: \[\frac{7}{\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}} = \frac{7\sqrt{2}}{2}\] Notice that in the denominator the square roots cancel out because the square root of a number times itself is that original number. So the answer is \[\frac{7\sqrt{2}}{2}\]